OpenStudy (anonymous):

Given \(\large \color{green}{ A = \frac{10}{\sqrt{2}}(a_x + a_z)}\) and \( \large \color{blue}{B = 3(a_y + a_z)}\), express the projection of \(B\) on \(A\) as a vector in the direction of \(A\)..

OpenStudy (anonymous):

Isn't it something like \[\text{proj}_AB=\frac{A\cdot B}{|A|}~~?\]

OpenStudy (anonymous):

Great, I thought I'd forgotten it :)

OpenStudy (anonymous):

But that will give you a scalar quantity at then end I think..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

IT is scalar product

OpenStudy (anonymous):

But Final answer is a vector in the book..

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

Sorry, but that was meaningless....

OpenStudy (anonymous):

Oh sorry, I think it's multiplied by \(A\). That makes a scalar multiple of the vector \(A\), and so it's in the same direction. \[\frac{A\cdot B}{|A|}A\]

OpenStudy (anonymous):

Hey, don't tag anybody, that is my work.. :P

OpenStudy (anonymous):

Yeah i know xD, I started vectors just a few days back

OpenStudy (anonymous):

Ganesh has seen it, may be he is trying on his first..

OpenStudy (anonymous):

You have multiplied vector A there???

OpenStudy (anonymous):

Okay I try that first.. A.B = 30/sqrt{2}

ganeshie8 (ganeshie8):

if you want vector along A, then multiply the mag by unit vector of A ?

OpenStudy (anonymous):

|A| = 10 Dividing we get 3/sqrt{2}

OpenStudy (anonymous):

And then multiplying by again vector A will give me: 30/sqrt{2}(a_x + a_z) and this is not the answer...

OpenStudy (anonymous):

Yeah, so you should have \[A\cdot B=\frac{30}{\sqrt2}a_xa_y+\frac{30}{\sqrt2}{a_z}^2\] \[|A|=\sqrt{\left(\frac{30}{\sqrt2}a_x\right)^2+\left(\frac{30}{\sqrt2}a_z\right)^2}=\frac{30}{\sqrt2}\sqrt{{a_x}^2+{a_z}^2}\] \[\frac{A\cdot B}{|A|}=\frac{a_xa_y+{a_z}^2}{\sqrt{{a_x}^2+{a_z}^2}}\] \[\frac{A\cdot B}{|A|}A={\large\frac{30\left(a_xa_y+{a_z}^2\right)}{\sqrt2\sqrt{{a_x}^2+{a_z}^2}}(a_x+a_z)}\]

OpenStudy (anonymous):

Wait Wait Wait.. How you got |A| that complex??

OpenStudy (anonymous):

I used the Pythagorean theorem. Isn't \(A\) given with variable coordinates?

OpenStudy (anonymous):

Oh sorry, May be you all did not get and I am sorry, I have not told you all earlier.. a_x, a_y and a_z are unit vectors along x, y and z direction..

OpenStudy (anonymous):

Oh well in that case just replace all \(a_x,a_y,a_z\) with 1.

OpenStudy (anonymous):

In A.B, a_x and a_y will become 0..

OpenStudy (anonymous):

\[A.B = (\frac{10}{\sqrt{2}} \cdot 3)a_z.a_z\]

OpenStudy (kainui):

You want B's magnitude in A's direction. So you divide out A's magnitude and find B's magnitude, and then multiply those together.

OpenStudy (anonymous):

\[A.B = \frac{30}{\sqrt{2}}\]

OpenStudy (anonymous):

Divide out A's magnitude with what??? @Kainui

OpenStudy (kainui):

With A's magnitude. \[\Large \frac{ \bar A}{| \bar A|}\] See, you just find the length of A, divide the vector A by the length and you will be left with a unit vector in the direction of A. It's essentially a directional form of the number 1 that you can multiply any scalar by to have this direction.

OpenStudy (anonymous):

Sorry, guys I will come somewhat later, my dinner is waiting for me.. Then I try it.. And @Kainui By that yes I will get unit vector along A's direction, then you want me to multiply it with B's magnitude???

OpenStudy (anonymous):

Okay, start over: \[A=\frac{10}{\sqrt2}(a_x+a_z)=\frac{10}{\sqrt2}\bigg(\langle 1,0,0\rangle+\langle0,0,1\rangle\bigg)=\frac{10}{\sqrt 2}\langle 1,0,1\rangle\\ B=3(a_y+a_z)=3\bigg(\langle 0,1,0\rangle+\langle0,0,1\rangle\bigg)=3\langle0,1,1\rangle\] Now, \[A\cdot B=0+0+\frac{30}{\sqrt2}=\frac{30}{\sqrt 2}\\ |A|=\sqrt{\left(\frac{10}{\sqrt2}\right)^2+0^2+\left(\frac{10}{\sqrt2}\right)^2}=\sqrt{\frac{200}{2}}=10\] \[\frac{A\cdot B}{|A|}A=\frac{3}{\sqrt2}A=\frac{30}{4}\langle1,0,1\rangle \]

OpenStudy (anonymous):

I am afraid, that is not matching with my final answer.. @SithsAndGiggles

OpenStudy (kainui):

@waterineyes exactly, that's all that's left to do. That's actually what the formulas are saying, it's just like condensed down and tends to leave out reasoning It's like we can think in terms of "square root of four" but the formula just says "2".

OpenStudy (anonymous):

That 4 should be a 2... So you can reduce to \[15\langle1,0,1\rangle\]

OpenStudy (kainui):

Sorry @SithsAndGiggles I didn't mean to be rude and come in.

OpenStudy (anonymous):

No worries! It's nice to have more people around to tell me if I made a mistake :)

OpenStudy (anonymous):

No, this way also I am not getting, @Kainui just solve it on your own and copy with my final answer, I am writing here...

OpenStudy (anonymous):

\[\frac{A}{|A|} = \frac{(a_x + a_z)}{\sqrt{2}}\]

OpenStudy (kainui):

Correct.

OpenStudy (anonymous):

\[\frac{A}{|A|} \cdot |B| = 3(a_x + a_z)\]

OpenStudy (kainui):

Yes. This is correct.

OpenStudy (kainui):

WAIT sorry.

OpenStudy (anonymous):

Why??

OpenStudy (kainui):

This is not correct. We are not looking for a vector of length B in the direction of A. X_X

OpenStudy (kainui):

 See the difference? The projection is the smaller one, the vector in the direction of A is much longer.

OpenStudy (anonymous):

Wait if I do: \[(B.a_A) \cdot a_A\]

OpenStudy (kainui):

This is still in the same direction as A, so A/|A| is still useful to us. But we can't multiply this by the magnitude of B. Instead the length projection is the dot product of B and A.

OpenStudy (anonymous):

We should again multiply it with the direction of A and direction of A is given by vector by its length..

OpenStudy (kainui):

Yes, the formula which Sith gave earlier is good, it just needs to be multiplied by the direction. Hahaha this is becoming an ordeal!

OpenStudy (anonymous):

In your terms If I do: \[(\frac{B. A}{|A|}) \cdot \frac{A}{|A|}\] I think I am getting answer by this way..

OpenStudy (kainui):

Yes this looks good now haha.

OpenStudy (anonymous):

Sith was saying to multiply with whole vector(Mag + Direc) but we have to multiply only with the direction...

OpenStudy (kainui):